# A Time-Intensity Aware Pipeline for Generating Late-Stage Breast DCE-MRI using Generative Adversarial Models

### v1

@@ Abstract
Contrast-enhancement pattern analysis is critical in breast magnetic resonance imaging (MRI) to distinguish benign from probably malignant tumors. However, contrast-enhanced image acquisitions are time-consuming and very expensive. As an alternative to physical acquisition, this paper proposes a comprehensive pipeline for the generation of accurate long-term (late) contrast-enhanced breast MRI from the early counterpart. The proposed strategy focuses on preserving the contrast agent pattern in the enhanced regions while maintaining visual properties in the entire synthesized images. To that end, a novel loss function that leverages the biological behavior of contrast agent (CA) in tissue, given by the Time-Intensity (TI) enhancement curve, is proposed to optimize a pixel-attention based generative model. In addition, unlike traditional normalization and standardization methods, we developed a new normalization strategy that maintains the contrast enhancement pattern across the image sequences at multiple timestamps. This ensures the prevalence of the CA pattern after image preprocessing, unlike conventional approaches. Furthermore, in order to objectively evaluate the clinical quality of the synthesized images, two metrics are also introduced to measure the differences between the TI curves of enhanced regions of the acquired and synthesized images. The experimental results showed that the proposed strategy generates images that significantly outperform diagnostic quality in contrast-enhanced regions while maintaining the spatial features of the entire image. This results suggest a potential use of synthetic late enhanced images generated via deep learning in clinical scenarios.
@@
A novel normalization strategy (TI-norm) for DCE-MRI imaging that preserves the kinetic properties of the TI curve across multiple sequences.
A novel loss function (TI-Loss) based on the time-intensity curve that ensures the prevalence of clinically relevant information in the generated images.
Two novel metrics based on the TI curve to objectively evaluate the performance of generated images in terms of clinically relevant information for interpretation.
@@ 1 Introduction
Early diagnosis and treatment of breast cancer can significantly improve effectiveness, leading to higher survival rates and reduced cancer-related deaths [1, 2]. Breast cancer is primarily detected through medical imaging, and Dynamic Contrast-Enhanced Magnetic Resonance Imaging (DCE-MRI) is considered the most sensitive method. DCE-MRI does not use radiation and allows the characterization of the internal structures of the tumor [3, 4]. Acquisition of DCE-MRI involves intravenously administration of a chemical gadolinium-based compound known as contrast agent (CA). Due to the accelerated metabolic activity of malignant tumors, they exhibit a characteristic CA absorption pattern that differs from benign and healthy tissues in both speed and intensity over time. This pattern is highly sensitive and makes tumor detection and characterization easier in comparison to other imaging modalities. For radiological findings (i.e. suspicious lesions), the CA absorption pattern is defined by the analysis of the kinetic or time-intensity curve (TI). The TI curve is computed from the variation of the pixel intensity in the location of the maximum intensity point from early or short-term (1-3 min) and late or long-term (8-10 min) responses. This analysis, along with visualization of the internal structure of tumors, makes DCE-MRI one of the most efficient tools for detecting and characterizing breast cancer. [5].
The acquisition of DCE-MRI varies even from hospital to hospital, as typically 5 to 8 screens are taken after CA administration, which usually takes around 15-20 minutes. However, no right number of shots or right time intervals have been established or standardized yet. Furthermore, the large number of images per patient requires a large infrastructure in terms of storage and transmission devices in the database center and the picture archiving and communication system (PACS). Consequently, radiologists must have a long interpretation time to reach a diagnostic conclusion, given the amount of information per patient to analyze. Thus, DCE-MRI becomes extremely restrictive in the number of patients that can be diagnosed; therefore, it is indicated only to individuals at higher risk.
@@
Alternatively, ultra-fast DCE-MRI (UF-DCE-MRI) has emerged to capture kinetic information in the very early post-contrast period by increasing the temporal resolution drastically, which significantly reduces acquisition time. However, increasing temporal resolution also decreases spatial resolution, leading to inaccuracies in the visualization of internal tissue structures [6]. This is why UF-DCE-MRI is not used today as a standalone examination. As another option, abbreviated MRI protocols have been proposed to reduce the number of image sequences that must be acquired, stored, and interpreted [7, 8]. Typically, these protocols include only one early post-contrast image sequence, which allows detection of image findings, but limits tumor characterization because they do not capture the dynamic TI pattern [9]. However, the late response is relevant to distinguish among enhanced masses with benign-appearing morphological features. For example, biologically aggressive cancers will exhibit strong angiogenic activity with fast wash-in and wash-out, whereas fibroadenomas often observe fast wash-in, although wash-out is scarce [10].
On the other hand, deep learning and machine learning models have shown the potential to generate synthetic screens from multiple medical imaging modalities, and several models have been proposed to generate missing or additional image sequences [11, 12, 13]. Some of them have been evaluated as an alternative for the reduction or elimination of CA doses [14, 15]. In the case of breast DCE-MRI, proposed models have focused mainly on synthesizing the early response from the pre-contrast image, which approximates the removal of CA administration [16]. To do so, adaptations of state-of-the-art models [17, 18] and the inclusion of optimization criteria based on the morphological characteristics of the images or tumors [19, 20] have been evaluated. In line with this trend, an intuitive path to obtain a full and useful virtual CA response for diagnosis will imply the synthesis of the late response. However, to our knowledge, no efforts have been reported to address the missing late CA response of breast DCE-MRI, besides our own presented in [21], where we introduced a temporal learning function to exploit the behavior of the TI curve to efficiently generate coherent DCE-MRI.
@@
Since the CA pattern of the TI curve is the most relevant information to distinguish benign from malignant lesions in DCE-MRI, the synthesis of contrast enhancement images must accurately simulate this pattern rather than synthesize a realistic image only. Taking this into account, this paper presents a comprehensive pipeline for the generation of long-term (late) contrast-enhanced breast MRI from its short-term (early) counterpart, which attempts to replicate the behavior of the TI curve of the enhanced regions while maintaining the visual properties in the entire image.
@@
The experimental results show that the proposed pipeline generates images that accurately synthesize the contrast enhancement response in terms of the TI curve for both radiological findings (regions of interest - ROI) and unannotated regions, while maintaining visual properties comparable to the real late contrast enhanced images. This result, achieved primarily by implementing the TI-Loss during generative model training, is extremely important, as the evaluation of the baselines indicates that optimizing for pixel-wise quality does not guarantee the prevalence of clinically relevant information. However, the use of TI-Loss alone does not achieve optimal results in terms of clinical performance. The role of the TI-norm was evaluated, demonstrating the prevalence of clinical interpretation and behavior among images to facilitate learning of their relationships while maintaining visual quality in the images. The comparative analysis of the models shows that the proposed pipeline allows the generation of images that preserve spatial properties in terms of pixel metrics, while significantly outperforming them in the clinical interpretation.
Our pipeline presents multiple contributions.
  • A novel normalization strategy (TI-norm) for DCE-MRI imaging that preserves the kinetic properties of the TI curve across multiple sequences.
  • A novel loss function (TI-Loss) based on the time-intensity curve that ensures the prevalence of clinically relevant information in the generated images.
  • Two novel metrics based on the TI curve to objectively evaluate the performance of generated images in terms of clinically relevant information for interpretation.
2 The Time-Intensity curve as diagnostic tool in DCE-MRI
DCE-MRI involves the use of a chemical gadolinium-based compound known as contrast agent (CA). The CA is injected intravenously before the MRI scans to distinguish certain tissue structures. This is possible because tumors or abnormal cells have faster metabolic activity, which makes CA absorb quickly. As a result of the active interaction between the gadolinium and the magnetic field of the MRI scanner, regions containing high concentrations of CA appear as luminous zones in the images. Due to the fast metabolic rate of tumor cells, CA enables precise visualization of the internal tumor composition, size, and shape. Furthermore, it is common to perform a temporal analysis of suspicious lesions across multiple MRI sequences acquired at specific time intervals to determine the most appropriate diagnosis and course of treatment for each patient. These benefits make this technique one of the most effective for cancer detection [5], and the entire collection of sequences and images is usually called DCE-MRI.
The variation pattern in pixel intensity for a zone in a DCE-MRI image over multiple time intervals is typically known as the kinetic or time-intensity (TI) curve. The TI curve is often used to determine the type of lesions based on the contrast enhancement pattern because it is easy to use and effective for both diagnosis and prognosis [23, 24, 25]. From the TI curve, lesions can be classified as probably benign or malignant based on the variation in CA absorption relative to normal tissue in the early and late phases [26]. As illustrated in Figure 1, the TI curve is typified according to its behavior as follows:
  • TI curve type I: This is defined as a progressive or increasing enhancement pattern. It is typical in the case of benign findings, where constant CA absorption but slow evacuation occurs. This increase is often greater than 10% w.r.t. early response over the long run.
  • TI curve type II: This is defined as a plateau pattern. This tendency implies early CA absorption that remains almost constant in the late stage and is equivocal about lesion benignity or malignancy. This variation is within 10% of the late response w.r.t. the early one.
  • TI curve type III: This is defined as a decreasing or washout pattern. This is typical in malignant tumors, where a significant increase in pixel intensities in the short time and a pronounced decrease in the longer time happen. Commonly, the late response decreases lower than 10% with respect to the early counterpart.
Figure 1: Illustration of different TI curves types for DCE-MRI. D0, D1, D2 and D3 stand for the dynamic sequences corresponding to weighted fat-saturated T1. D0 corresponds to pre-contrast, D1 to early post-contrast and D2-D3 to late post-contrast. Blue dotted line shows a persistent TI curve (enhancement increase higher than +10%), orange dashed line shows a plateau TI curve (enhancement change between +10% and -10%) and red dot-dashed line shows a wach-out TI curve (enhancement decrease lower than -10%).
The time-intensity curve has been reported to be an important tool for the clinical evaluation of radiological findings or abnormalities in the case of multimodal DCE MRI. This provides a solid evaluation framework from the pixel analysis level, with reproducible results [27, 28]. The TI curve is widely used in multiple scenarios, as it provides diagnostic value to differentiate benign lesions[29, 30], as well as their type of characterization [31]. In addition, the time intensity curve has been explored and reported as a powerful tool for multiple purposes. To predict the response to treatment in patients [32, 33]. As a key tool for improving and boosting the diagnosis of ultra-fast DCE-MRI by leveraging the increase in temporal resolution [34, 35]. Finally, as an assessment tool to determine the validity of low- or reduced-CA-dose approaches in DCE-MRI [36, 37].
@@ 3 Methods
A graphical representation of the proposed pipeline is presented in Figure 2. It is composed by three main stages: Generative model training, late post-contrast synthesis and model evaluation. Before training the generative model, training images are normalized by using a novel normalization technique named time-intensity normalization (TI-norm), that attempts to avoid losing the intensity changes among dynamic sequence, i.e. pre-contrast ant post-contrast images. Then, a Pixel Attention Network (PAN) is used as backbone to synthesize the late post-contrast image from the early counterpart. A novel loss function called TI-loss is here proposed to guide the learning process to minimize the difference among the real and late CA responses. The trained PAN using the TI-loss (TI-PAN) is then used for generating a synthetic late post-contrast image from the TI normalized early image. Finally, the performance of the late contrast synthesis is twofold evaluated. First, conventional image quality metrics (MAE, SSIM and PSNR) between real and generated late post-contrast images are computed to determine the spatial quality of the generated images. Second, two metrics to determine the difference between the real and generated time-intensity curves, allowing the evaluation of the ability to maintain the clinical significance of the CA pattern in the synthesized image.
@@ 3.1 Time Intensity Pixel Attention Network: TI-PAN
As mentioned above, the proposed pipeline is based on synthesizing in the best way the CA pattern. We used an attention-based model called Pixel attention network (PAN) [38]. This model uses attention layers under unique scheme blocks known as self-calibrated convolution (SC-PA) for feature extraction and nonlinear mapping, and nearest-neighbor upsampling (U-PA) for reconstruction. These blocks are considerably more effective than their conventional residual/dense counterparts at the lowest parameter cost. Once the model receives an early-response input image $x_{e}$, the PAN extracts features at the spatial level and stacks multiple SC-PA blocks to generate powerful representations. Subsequently, U-PA blocks are used as reconstruction modules to upsample the features. After reconstruction, a global skip connection path is used in which a bilinear interpolation to the input is performed. Then, the output corresponding to the late response $y_{l^{\prime}}$ is obtained. Finally, to encourage the realism in the image, we added a discriminator network and included an adversarial approach such as in generative adversarial networks (GANs) 3. This also included the pixel-wise loss based on the $L1$ norm to enforce spatial dependency. The full objective for the PAN model is shown in equation 9
@@
To help generative models understand contrast agent behavior in tissue, we aim to formulate a general function that models the behavior of the CA response in the images. To achieve this, we created a loss term to determine the difference in the enhanced areas by the contrast agent called the Contrast Enhancement Loss (CELoss) [21]. Specifically, we consider the real early ($x_{e}$) and late ($y_{l}$) responses sampled from data $X_{e}$ and $Y_{l}$ respectively; we compute the real contrast enhanced map ($CE_{r}(x_{e},y_{l})$) by subtracting early from late response (equation 5) using the $L1$ norm. In a similar sense, the contrast enhanced map is also computed for the generated image $CE_{g}$ (equation 6). The model is trained to minimize the difference between $CE_{r}$ and $CE_{g}$ using the $L1$ norm as formulated in equation 7.
@@
Hence, combining the TI-Loss with the adversarial approach of the PAN could be defined as in equation 9, where $\lambda$ is a weighting hyperparameter for the loss function.

### v2

# Synthesizing Late-Stage Contrast Enhancement in Breast MRI: A Comprehensive Pipeline Leveraging Temporal Contrast Enhancement Dynamics

@@ Abstract
Dynamic contrast-enhanced magnetic resonance imaging (DCE-MRI) is essential for breast cancer diagnosis due to its ability to characterize tissue based on contrast agent kinetics. conventional DCE-MRI protocols require multiple imaging phases, including both early and late post-contrast acquisitions, leading to prolonged scanning times that can cause patient discomfort and motion artifacts, as well as contribute to higher costs and limited availability in clinical settings. To address these limitations, this paper presents a comprehensive pipeline for synthesizing long-term (late-phase) contrast-enhanced breast MRI images from short-term (early-phase) counterparts, aiming to replicate the behavior of the time-intensity (TI) curve in enhanced regions while maintaining visual properties across the entire image. The proposed approach introduces a new loss function called the Time Intensity Loss (TI-loss), which leverages the temporal behavior of the contrast agent to guide the training of a generative model. Furthermore, as established normalization strategies shown undiserable effects on the enhancement beahviour, a novel normalization strategy (TI-norm) is also proposed, which preserve the contrast enhancement pattern across multiple image sequences at various timestamp. Additionally, two new metrics are proposed to evaluate the synthesized image quality, i.e., the Contrast Agent pattern score ($\mathcal{CP}_{s}$), which determines the validity of annotated regions according to their enhancement patterns (plateau, persistent, washout), and the average difference in enhancement ($\mathcal{ED}$), that quantify the difference between the real and generated enhancement in selected regions. Evaluation was performed using public DCE-MRI dataset that includes studies from 3T and 1.5T scanners with different imaging techniques. Experimental results demonstrate that our method accurately synthesizes the contrast enhancement response in terms of the TI curve in regions of interest, that significantly outperforms other models, while maintaining visual properties comparable to real late-phase contrast-enhanced images. By enabling accurate synthesis of late-phase contrast-enhanced images from early-phase data, our method has the potential to optimize DCE-MRI protocols, reducing scanning time without compromising diagnostic accuracy. This advancement brings generative models closer to practical implementation in clinical scenarios, enhancing efficiency in breast cancer imaging.
@@
Title: Synthesizing Late-Stage Contrast Enhancement in Breast MRI: A Comprehensive Pipeline Leveraging Temporal Contrast Enhancement Dynamics
  • Ruben D. Fonnegra*
    Affiliation: Institución Universitaria Pascual Bravo, Instituto Tecnológico Metropolitano
    City/Country: Medellín, Colombia
    Email: ruben.fonnegra@pascualbravo.edu.co
    Permanent email address: rubenfonnegrat1015@gmail.com
  • Maria Liliana Hernández
    Affiliation: Ayudas Diagnósticas SURA S.A.S.
    City/Country: Medellín, Colombia
  • Juan C. Caicedo
    Affiliation: Morgridge Institute for Research, Dept. Biostatistics and Medical Informatics, University of Wisconsin—Madison
    City/Country: Madison, WI, USA
  • Gloria M. Díaz
    Affiliation: Instituto Tecnológico Metropolitano
    City/Country: Medellín, Colombia
@@
A comprehensive pipeline for synthesizing late-phase contrast-enhanced breast MRI images from early-phase that ensures coherent image synthesis and accurate clinical interpretation, bringing generative models closer to practical implementation in clinical scenarios.
A new loss function (TI-Loss) that leverages the temporal behavior of contrast agents to guide the training of generative models, ensuring accurate replication of contrast enhancement patterns.
A novel normalization strategy (TI-norm) for DCE-MRI imaging that preserves the kinetic integrity of the time-intensity curve across multiple image sequences and timestamps showing a critical effect during model training.
Two new TI curve-based metrics to objectively assess the quality and clinical reliability of generated images according to their real biological behavior.
@@ 1 Introduction
Breast cancer remains one of the leading causes of morbidity and mortality among women worldwide, and early detection and accurate diagnosis are critical for improving prognosis and reducing the mortality rate associated with it  [1, 2]. Dynamic contrast-enhanced magnetic resonance imaging (DCE-MRI) has become an essential tool in evaluating breast lesions due to its high sensitivity and capability to characterize breast tissue and the internal structures of the tumor  [3, 4]. This is made possible by the intravenous administration of a gadolinium-based compound known as a contrast agent (CA). Due to the accelerated metabolic activity of malignant tumors, they exhibit a characteristic CA absorption pattern over time that differs from benign and healthy tissues in both speed and intensity. This pattern is highly sensitive and facilitates tumor detection and characterization, making it easier in comparison to other imaging modalities. For radiological findings (i.e. suspicious lesions), the CA absorption pattern is defined by the analysis of a curve known as the kinetic or time-intensity curve (TI), which is computed from the variation of pixel intensity at the location of the maximum intensity point from early or short-term (1-3 min) and late or long-term (8-10 min) responses. This analysis, along with the visualization of the internal structure of tumors, makes DCE-MRI one of the most efficient tools for detecting and characterizing breast cancer  [5].
@@
Although DCE-MRI protocols vary even from hospital to hospital, conventionally, they involve acquiring multiple sequences after CA administration, including at least both early and late contrast response images, which significantly prolongs the scanning time. This extended duration can cause discomfort for patients, increase the risk of motion artifacts, and limit the availability of MRI equipment for other examinations. Furthermore, the large number of images per patient requires substantial infrastructure in terms of storage and transmission devices in the database center and the picture archiving and communication system (PACS). Consequently, radiologists must spend a considerable amount of time interpreting images to reach a diagnostic conclusion, given the volume of information per patient to analyze. Alternatively, ultra-fast DCE-MRI (UF-DCE-MRI) has emerged to capture kinetic information in the very early post-contrast period by drastically increasing the temporal resolution, which significantly reduces acquisition time. However, increasing temporal resolution also decreases spatial resolution, leading to inaccuracies in the visualization of internal tissue structures  [6]. As another option, abbreviated MRI protocols have been proposed to reduce the number of image sequences that must be acquired, stored, and interpreted  [7, 8]. Typically, these protocols include only one early post-contrast image sequence, which allows for the detection of image findings or lesions but limits tumor characterization because they do not capture the dynamic time-intensity (TI) pattern that distinguishes between probably benign and probably malignant enhanced masses  [9]. As an example, biologically aggressive cancers will exhibit strong angiogenic activity with fast wash-in and wash-out, whereas fibroadenomas often exhibit fast wash-in, although wash-out is scarce  [10].
Recently, generative artificial intelligence has shown the potential to synthesize images, and numerous studies have demonstrated the feasibility of using these models in several medical imaging domains [11, 12, 13]. Particularly, synthesis of contrast-enhanced images has been proposed as an alternative to reducing or eliminating the need for contrast agents [14, 15]. In the case of breast DCE-MRI, sSome studies have explored generating virtual contrast images from non-contrast MRI scans, aiming to eliminate the need for contrast agents altogether [16]. Consequently, generative models incorporating specific pixel-based metrics or morphological features have been presented to surpass precise local tumor visualization [17, 18, 19, 20]. Nonetheless, these works have been constrained by their limited ability to furnish relevant information that extends beyond their realism for diagnostic purposes [21].
Initial studies focused on generating the first post-contrast sequence from pre-contrast or non-contrasted images through the use of generative artificial intelligence methods  [17, 18, 22]. Furthermore, to enhance image quality in the visualization domain, subsequent studies have suggested incorporating additional visual cues, such as morphological or shape features.  [19, 20, 23]. Simultaneously, other authors have assessed the utilization of different image sequences, including diffusion-weighted and apparent diffusion coefficient images, regarding their influence on the visual domain of synthesized images [24, 25], and T2-weighted images [26, 27]. Recently, a comprehensive evaluation of several input sequences for generating virtual contrast-enhanced images showed that, even though quantitative metrics indicate similar performance in almost all cases, there are significant differences in qualitative evaluations regarding qualitative scores, including diagnostic image quality, image sharpness, satisfaction with image contrast, and visual signal-to-noise ratio  [21]. Thus, despite the fact that Although these methods hold significant technological promise, their current lack of clinical confidence and inability to capture essential contrast dynamics necessary for accurate lesion characterization cast doubt on their near-term applicability in clinical practice.
In order to address the diagnostic need for temporal evaluation of lesion dynamics and obtain a comprehensive and useful virtual CA response for diagnosis, a few works have proposed synthesizing the late-phase contrast image. Oh et al. [28] proposed generating a ”late-phase” (210s after CA administration) from the early UF-DCE-MRI sequence (70 and 140s after CA administration) using a modified paired StarGAN model and the difference phase map to specifically focus on tumor information during training. The approach’s results are promising; however, it is based on UF-DCE-MRI, which has low spatial resolution. Furthermore, generated images are equivalent to early post-contrast in conventional DCE-MRI and may not capture the required long-term information. Li et al. [29] introduced a multi-temporal fusion network that fuses low-dimensional attention maps from the first and third (early) post-contrast phase images to generate the eighth phase image (late). Feature fusion is performed using a pixel-level information co-attention model that attempts to highlight enhancement regions where a contrast agent response could be found. While this study has effectively used temporal image sequences to predict late responses, it still relies on a pixel-only technique, which could fail to provide assessments that are both interpretable and clinically pertinent. This gap underscores the need for strategies and evaluation frameworks that align with the diagnostic tools and interpretative practices widely accepted in the radiological community.
Conversely, we propose attempts to capture the dynamic behavior of the contrast agent and use this to train generative model frameworks that synthesize late-phase images, preserving this behavior while maintaining spatial characteristics throughout the entire image. In our prior research [30], we presented a loss function that approximates the expected reaction response of the contrast agent in tissue by reducing the overall difference between the real and predicted enhancement patterns. The results of this study motivated the exploration of generative models aimed at emulating the clinical information response within a generative imaging framework.
@@
Since the CA pattern of the TI curve is the most relevant information to distinguish benign from malignant lesions in DCE-MRI, the synthesis of contrast enhancement images must accurately simulate this pattern rather than synthesize a realistic image only. With this aim, this study presents a robust and comprehensive pipeline for creating long-term (late) contrast-enhanced breast MRI images from their short-term (early) counterparts. This method effectively models the TI curve behavior in enhanced regions while maintaining visual and structural integrity throughout the image, thereby addressing significant deficiencies in clinical imaging workflows and establishing a new standard for the production of high-quality, diagnostically pertinent MRI outputs. This pipeline intends to synthesize images that preserve accurate interpretability to approach generative models for their implementation in clinical scenarios.
@@
The experimental results show that the proposed pipeline generates images that accurately synthesize the contrast enhancement response in terms of the TI curve for both radiological findings (regions of interest - ROI) and unannotated regions, while maintaining visual properties comparable to the real late contrast-enhanced images. This result, achieved primarily by implementing the TI-Loss during generative model training, is extremely important, as the evaluation of the pixel-only models indicates that optimizing for pixel-wise quality does not guarantee the prevalence of clinically relevant information. However, the use of TI-Loss alone does not achieve optimal results in terms of clinical performance. The role of the TI-norm was evaluated, demonstrating the prevalence of clinical interpretation and behavior among images to facilitate learning of their relationships while maintaining visual quality in the images. The comparative analysis of the models shows that the proposed pipeline allows for the generation of images that preserve spatial properties in terms of pixel metrics, while significantly outperforming them in clinical interpretation.
@@ 2 Methods
A graphical representation of the proposed pipeline is presented in Figure 1. It is composed of three main stages: generative model training, late post-contrast synthesis, and model evaluation. Before training the generative model, training images are normalized using a novel normalization technique named time-intensity normalization (TI-norm), which attempts to avoid losing the intensity changes among dynamic sequences, i.e. pre-contrast and post-contrast images. Then, a Pixel Attention Network (PAN) is used as the backbone to synthesize the late post-contrast image from the early counterpart. A novel loss function called TI-loss is proposed here to guide the learning process to minimize the difference between the real and late CA responses. The trained PAN using the TI-loss (TI-PAN) is then used to generate a synthetic late post-contrast image from the TI-normalized early image. Finally, the performance of the late contrast synthesis is evaluated in twofold. First, conventional image quality metrics (MAE, SSIM and PSNR) between real and generated late post-contrast images are computed to determine the spatial quality of the generated images. Second, two metrics are used to determine the difference between the real and generated time-intensity curves, allowing for the evaluation of the ability to maintain the clinical significance of the CA pattern in the synthesized image.
@@ 2.1 Time Intensity Pixel Attention Network: TI-PAN
As mentioned above, the proposed pipeline is based on synthesizing the CA pattern in the best way. We used an attention-based model called the Pixel AAttention Ntwork (PAN) [32]. This model employs attention layers under a unique scheme of blocks known as self-calibrated convolution (SC-PA) for feature extraction and nonlinear mapping, and nearest-neighbor upsampling (U-PA) for reconstruction. These blocks are considerably more effective than their conventional residual/dense counterparts at a lower parameter cost. Once the model receives an early-response input image $x_{e}$, the PAN extracts features at the spatial level and stacks multiple SC-PA blocks to generate powerful representations. Subsequently, U-PA blocks are used as reconstruction modules to upsample the features. After reconstruction, a global skip connection path is employed, wherein a bilinear interpolation to the input is performed. Then, the output corresponding to the late response $y_{l^{\prime}}$ is obtained. Finally, to encourage realism in the image, we added a discriminator network and incorporated an adversarial approach, similar to that used in generative adversarial networks (GANs) 3. This also includes the pixel-wise loss based on the $L1$ norm to enforce spatial dependency. The full objective for the PAN model is shown in equation 9
@@
To help generative models understand contrast agent behavior in tissue, we aim to formulate a general function that models the behavior of the CA response in the images. To achieve this, we created a loss term to determine the difference in the enhanced areas by the contrast agent, called the Contrast Enhancement Loss (CELoss)  [30]. Specifically, we consider the real early ($x_{e}$) and late ($y_{l}$) responses sampled from data $X_{e}$ and $Y_{l}$ respectively; we compute the real contrast-enhanced map ($CE_{r}(x_{e},y_{l})$) by subtracting the early response from the late response (equation 5) using the $L1$ norm. In a similar manner, the contrast-enhanced map is also computed for the generated image $CE_{g}$ (equation 6). The model is trained to minimize the difference between $CE_{r}$ and $CE_{g}$ using the $L1$ norm as formulated in equation 7.
@@
Hence, combining the TI-Loss with the adversarial approach of the PAN can be defined as shown in equation 9, where $\lambda$ is a weighting hyperparameter for the loss function.